### v1

# Studying the Effects of Collaboration in Interactive Theme Discovery Systems

Abstract
NLP-assisted solutions have gained considerable traction to support qualitative data analysis. However, there does not exist a unified evaluation framework that can account for the many different settings in which qualitative researchers may employ them. In this paper, we take a first step in this direction by proposing an evaluation framework to study the way in which different tools may result in different outcomes depending on the collaboration strategy employed. Specifically, we study the impact of synchronous vs. asynchronous collaboration using two different NLP-assisted qualitative research tools and present a comprehensive analysis of significant differences in the consistency, cohesiveness, and correctness of their outputs.
11footnotetext: Equal Contribution
1 Introduction
Making sense of large amounts of textual data is a common task in many academic disciplines and is generally achieved through qualitative methods such as Thematic Analysis (Braun and Clarke, 2006) and Grounded Theory (Glaser et al., 1968). In these methods, researchers manually search through documents to identify common abstract themes that emerge from the data in a process commonly referred to as "inductive coding". However, manual inductive coding becomes untenable as datasets increase exponentially in size, leading researchers to adopt Natural Language Processing (NLP) solutions to automate parts of the coding process Brady (2019); Hilbert et al. (2019).
In the last decade, several NLP solutions have been proposed to support qualitative research methods. These systems offer different ways to automatically wrangle large amounts of textual data to identify common patterns or themes, such as discovering hidden semantic structures using topic models Smith et al. (2018); Fang et al. (2023), grouping documents based on similarity and broadcasting small sets of human annotations to the entire dataset Pacheco et al. (2023), or providing live recommendations to users as they manually code the data Dai et al. (2023); Gao et al. (2024). To ensure the agency of the researchers, all of these systems seek to strike a balance between automation and manual coding by incorporating some type of human-in-the-loop (HitL) strategy.
In most cases, these tools are evaluated to highlight the strengths and weaknesses of the proposed system. Examples include evaluating topic coherence with and without human input Fang et al. (2023) and contrasting machine-assisted code-book generation with a manual-only code-book generation Dai et al. (2023). In this paper, we argue that these evaluation strategies are often too narrow and do not account for the different real-world settings in which qualitative researchers employ HitL systems. In particular, we want to study how tools for inductive coding may result in different outcomes in different collaborative settings.
Collaborative data analysis occurs when two or more researchers pool their focus into a shared body of data to produce an agreed interpretation. Potential advantages of collaboration include producing more reliable results, a richer understanding of the data, and knowledge that transcends individual perspectives Flick (2014). While there are many ways in which collaboration can occur, in this work we look at two common but opposite qualitative research collaboration settings. The first setting is defined by individual coders working separately before jointly consolidating their findings (asynchronous coding). The second setting is defined by coders working in groups and identifying themes through synchronous discussion (synchronous coding). Each setting has different strengths: asynchronous coding can be a more flexible option for teams as they can work across different regions and time zones, whereas synchronous coding can combine diverse perspectives and streamline the annotation processes.
We test these settings by choosing two NLP tools for inductive coding, and performing coding experiments in the synchronous and asynchronous contexts. We choose tools that use different underlying NLP methods for this work: (1) the Fang et al. (2023) system, which uses a traditional human-in-the-loop topic modeling solution, and (2) the Pacheco et al. (2023) system, which models emerging themes with respect to their relationship to other high-level concepts, groups similar instances based on this, and leaves theme definitions and annotations to users, which are then propagated to the full dataset. This way, we can have a more representative understanding of how different tools can be affected in synchronous and asynchronous collaboration.
To measure outcomes in synchronous vs. asynchronous collaboration, we propose a comprehensive evaluation strategy that looks at different output qualities. Namely, we are interested in measuring the consistency of outputs between synchronous and asynchronous coding, as well as the cohesiveness, and correctness of the resulting themes within each collaboration setting. Finally, we also perform a small-scale user study to draw insights from the experience of the coders under each collaboration setting.
In summary, we make two main contributions:
  1. 1.
    We demonstrate how the collaboration setting in which qualitative research takes place could affect the outcome of solutions proposed in the NLP literature, by testing two NLP-assisted inductive coding tools in synchronous and asynchronous coding contexts.
  2. 2.
    We propose a comprehensive evaluation strategy that measures different aspects of the outcome of NLP-assisted inductive coding tools, offering insights into the types and quality of themes that they are able to capture.
We hope that our findings will inform the future development of language technologies that are in line with the expectations and workflow of qualitative researchers, as well offer insights into how the next generation of qualitative research tools should be evaluated so that researchers are able to choose the right tool for their coding requirements.
2 Related Work
Our work builds on previous literature on computational solutions for qualitative coding. The overarching goal of these systems is to partially automate the qualitative coding process by either learning user-defined themes interactively Pacheco et al. (2023); Gao et al. (2023) or by inducing topics or clusters in an interactive, semi-supervised manner Fang et al. (2023); Smith et al. (2018). A separate but related line of work exemplified by Gao et al. (2024) uses language models to generate label recommendations as the user performs the coding process. While this system is explicitly designed for asynchronous collaboration, the systems we study differ in their ability to classify large portions of the dataset without requiring extensive supervision.
Our research addresses a real-world use case for qualitative researchers using HiTL systems and is informed by literature from the field of Human-Computer Interaction (HCI)  Jiang et al. (2021); Feuston and Brubaker (2021); Chen et al. (2018). HCI research centers the user as the focus of study in order to identify where and how best to place NLP tools in interactive systems. We build on these findings by proposing a comprehensive method for evaluating a given system for qualitative coding in addition to conducting user interviews. To this end, we also center part of our experiment on the user to uncover differences in the interactive experience between synchronous and asynchronous environments.
A common finding in the HCI studies surveyed is the researchers’ focus on ambiguity. In traditional settings, researchers can independently code a significant portion of the dataset and meaningfully identify areas of ambiguity. However, when working with large datasets and relying on NLP systems, researchers cannot easily uncover all such areas. Solutions have been proposed to either visualize codes (Drouhard et al., 2017) or rank document disagreement (Zade et al., 2018) regardless of dataset size. Our evaluation differs from these works since we look at different output qualities of the resulting themes and their assignments, combining signal from group overlaps, relationships in the semantic embedding space, as well as manual post-hoc evaluations. This method of evaluation highlights areas where the coders diverge both with each other and with the model, providing another perspective on the ambiguity question.
The motivating factor for our study is the popularization of HitL systems for qualitative coding in the NLP field. With the growing number of new interactive systems comes a need for standardized evaluation across different real-world environments. Previous evaluation methods introduced with new systems were generally ad hoc, with experiments conducted variously in group settings  Choo et al. (2013); Hoque and Carenini (2016); Smith et al. (2018), on individual participants  Rietz et al. (2020), and through platforms such as MTurk  Zade et al. (2018). Our contribution provides a standardized framework for performing experiments in synchronous and asynchronous settings, as well as a set of metrics for evaluating consistency, cohesiveness, and correctness in experimental results.
3 Interactive Systems
We evaluate two interactive theme discovery systems that use different approaches to accomplish the same task. This way, we ensure that our analysis can account for variation in NLP techniques and annotation interfaces.
3.1 Fang et al. (2023)
This system is a HitL, query-driven topic model (QDTM) (Fang et al., 2021) that allow researchers to iteratively adjust clusters using a set of refinement operations. Broadly, the QDTM is a semi-supervised topic modeling algorithm that accepts input queries to extract relevant concept words. The extracted concept words are used in a two-phase framework based off the Hierarchical Dirichlet Process (HDP) (Teh et al., 2004) to perform topic modeling. In the first phase, the HDP is used to infer one "parent topic" for each concept along with irrelevant topics, where each topic is defined by a set of weighted words. In the second phase, the model expands the parent topic into a set of subtopics using the Gibbs sampling process. The words for each subtopic are added to the parent topic to produce a final set of topics, which are used to assign a distribution of topics to each document.
Given an initial topic model, users can add, remove, or swap the order of words within each topic, which alters the probability of topic assignment for documents that contain that specified word. Similarly, users can split or merge topics based on topic words; doing so does not guarantee that assigned documents will be reassigned accordingly due to the stochastic nature of the model. Users can also remove documents from a given topic, but cannot directly add documents in the same manner. In each iteration, users may freely select refinement operations for each topic before applying all selections at once. Users can undo selections prior to applying refinements and previous models are also saved if users wish to return to a previous iteration.
The overarching goal of this system is to facilitate the discovery of an optimal topic model for a given input dataset. By applying a selection of refinements, users can compare the effects of different topic word allocations on the resulting model. Returning to previous models and applying new refinements creates a branching model, which provides users with a wide space for exploration. Once satisfied with a given iteration, the user can download that model with its document distribution.
3.2 Pacheco et al. (2023)
This system uses a two-stage framework to assist researchers in characterizing large textual collections. In the first stage, the system automatically proposes an initial partition of the data, such that textual instances that are thematically similar are clustered together. Users are then provided with a set of exploratory operations that allows them to evaluate the quality of the discovered partitions, as well as to further explore and partition the space by inspecting individual examples, finding similar instances, and using open text queries. Next, users are prompted to group identified patterns into general themes, and instantiate them by giving them a name. Once a theme is created, users are provided with a set of intervention operations to explain the themes using natural language, select good example instances, write down additional examples, and input or correct supporting concepts to characterize the theme assignments. The full set of operations supported by the system are outlined in App. A.
In the second stage, the system finds a mapping between the full set of instances and the themes instantiated by the users. It takes the information contributed by the users in the form of examples and supporting concepts, and learns to map instances to themes using a neuro-symbolic procedure implemented with DRaiL Pacheco and Goldwasser (2021). The main goal of this process is to condition new theme assignments not only on the embedding distance between instances and provided examples, but also leverage the additional judgements provided by users via supporting concepts. For example, when analyzing a corpus about the Covid-19 vaccine, users could point out that 80% of the examples for theme Government Distrust have a clear anti-vaccine stance. The system then uses this information to introduce inductive bias into the mapping procedure. DRaiL uses weighted first-order logic rules to express decisions and dependencies between different decisions, for example $w:\mathtt{stance(anti\_vax)}\rightarrow\mathtt{theme(gov\_distrust)}$. For more details about this implementation, we refer the reader to the original paper.
During this process, the systems keeps instances unassigned to a theme if there is not a good enough match. Then, a consecutive partitioning step is done over all unassigned instances. Finally, users have the option to repeat this two-stage process until convergence.
4 Coding Protocol
In order to study the performance of the two systems discussed in Section 3 under different collaboration settings, we design a protocol to conduct qualitative coding in the synchronous and asynchronous qualitative coding paradigms. We use the outcome of our coding described here to gather results, and then conduct an evaluation in the subsequent sections.
4.1 Participants
We recruited a group of 15 researchers in NLP and Computational Social Science, 3 female and 12 male, within the ages of 20 and 45. This group included professors at different levels of seniority, postdoctoral researchers, and graduate and undergraduate students from two different universities. All participants were either well versed in qualitative data analysis, or were explicitly trained by more senior researchers to perform the task.
4.2 Dataset
We use a common dataset to perform qualitative coding using the Pacheco et al. (2023) and Fang et al. (2023) systems. This dataset was originally released by Pacheco et al. (2022), and consists of approximately $85,000$ Tweets about the COVID-19 vaccine. All tweets in this corpus were posted by users located in the United States, are uniformly distributed between Jan. and Oct. 2021, and contain predictions for vaccination stance (e.g. pro-vax, anti-vax) and morality frames (e.g. fairness/cheating and their actor/targets.) Haidt and Graham (2007); Roy et al. (2021).
4.3 Qualitative Coding Protocol
In order to compare the consistency between different collaborative settings, we need at least two sets of synchronous/asynchronous annotations per annotation system. To ensure compatibility between synchronous and asynchronous outcomes, we use the same starting clusters for each system. The hyperparameters used in the Fang et al. (2023) system are the same as in their study and we did not use any input queries prior to initialization. We scheduled annotations for each setup separately, with details outlined in the sections below.
4.3.1 Synchronous Coding
Coding for the Pacheco et al. (2023) system was done over Zoom, where participants met online and collaboratively discussed and annotated the dataset. Six participants were bifurcated into two groups of three individuals, with each group communicating only with their members during the annotation process. Coders followed the protocol outlined in the original paper, performing two consecutive rounds of interaction, each round lasting a total of 1 hour.
For the Fang et al. (2023) system, we selected a different set of six annotators and divided them into two separate groups of three individuals each. Annotators met with their group and collaboratively coded the data. Prior to the annotation session, participants were briefed about the different features of the system and how to use them. The annotators went over the pre-assigned topics proposed by the system, and modified them to make them more accurate and cohesive. Both groups of annotators reached saturation after multiple passes over the data, in sessions lasting around 2.5 hours.
4.3.2 Asynchronous Coding
The asynchronous coding processed followed the same outline as the synchronous coding for each system, except that it was conducted by three different annotators, and it was done individually without any collaboration. For each system, the annotators first met in-person and started annotations separately. The purpose of the initial co-located coding session was to troubleshoot any problems and solve any doubts that the participants may have. The participants then completed the annotation completely asynchronously and remotely.
Pacheco et al. (2023) Fang et al. (2023)
Jaccard Sim. Centroid Sim. Group Avg. Sim. Jaccard Sim. Centroid Sim. Group Avg. Sim.
Sync $\bf 0.36(0.19)$ $\bf 0.98(0.01)^{*}$ $\bf 0.52(0.07)^{*}$ $\bf 0.56(0.23)$ $\bf 0.98(0.05)^{**}$ $\bf 0.52(0.10)$
Asyn $0.30(0.22)$ $0.94(0.07)^{*}$ $0.44(0.10)^{*}$ $0.30(0.17)$ $0.96(0.05)^{**}$ $0.51(0.09)$
Table 1: Avg. Consistency between Best Theme Matches across Annotator Groups. Statistically significant using a two-sample unpaired t-test with $p<0.05$. ∗∗ Near statistically significant with $p\approx 0.05$.
5 Evaluation
In this section, we use both descriptive metrics and a user study to provide a comprehensive analysis on the differences when annotating in synchronous and asynchronous settings. We first define and present several metrics for theme induction quality based on overlap, similarity/distance, association and verification that are well-established in the literature Ben-David and Ackerman (2008); Hoyle et al. (2021); Pacheco et al. (2023). These metrics are calculated using the sets of documents assigned to each theme by their respective system. While the Pacheco et al. (2023) system explicitly assigns a single theme for each document, the Fang et al. (2023) system provides a distribution of theme weights for each document. To ensure that results from both systems are comparable, we assign each document to the theme with the highest weight.
Finally, we summarize annotator experiences to provide a perspective on qualitative differences between the two settings. Combining dimensions in our evaluation methodology, we demonstrate that there are significant differences in consistency, cohesiveness, correctness, and user experiences between synchronous and asynchronous experiments.
5.1 Consistency
Coding individually always runs the risk that coders overgeneralize or miss key themes, resulting in “unconstrained or unsystematic” analysis (Cornish et al., 2014). We propose consistency as one dimension for measuring this systematicity. Building on the idea of inter-coder reliability, we define consistency as the measure of annotators eliciting the same themes from the same set of texts.
There are two obstacles to measuring annotation consistency when using semi-automated annotations. First, themes with similar names may not share the same annotated documents and themes with different names may share documents that are semantically similar, which we address by measuring the consistency of documents between given pairs of themes. Following the definition of consistency in Section 4 of Pacheco et al. (2023), we calculate the Jaccard similarity given as:
$J(A,B)=\frac{|A\cap B|}{|A|+|B|-|A\cap B|}$ (1)
where $A$ and $B$ are sets of documents for two themes with different coders. We then select the maximum Jaccard similarity for each (unique) theme, which we consider to be the “best theme match”, and calculate their averages for comparison across settings.
Second, since the majority of documents are assigned a theme by the system, the same themes may contain a different assortment of documents but remain relatively semantically similar. We therefore also measure the semantic similarity between themes using their SBERT embeddings Reimers and Gurevych (2019). We do this in two ways: (1) Centroid similarity, where the similarity of two themes is the similarity between their two centroids, and centroids are calculated by taking the average of all document embeddings within the theme. (2) Group avg. similarity, where the similarity between two themes is the average of pair-wise similarity between all documents in the two themes. As in the Jaccard case, we report the maximum embedding similarity for each (unique) theme, which we consider to be the “best theme match”, and calculate their averages for comparison across settings. In all cases, we use the cosine similarity.
Table 1 shows the average maximum Jaccard and embedding similarities for each experiment. Comparing the results on the synchronous/asynchronous dimension, we find that synchronous groups produce themes with higher average consistency across the board. The centroid similarities and group average similarities in particular are statistically significant, supporting our hypothesis that different settings produce noticeably different coding results. These findings also demonstrate that the theoretical advantage for collaborative coding holds empirically true when using HitL systems as well: coders are more likely to miss themes or overgeneralize when working independently.
Comparing between consistency metrics, we find that centroid similarity is at or near statistical significance for all experiments and is much higher than the other metrics across the board. This finding is also in line with our expectations detailed above: since HitL systems automate a large portion of document assignments, small differences in user interactions can result in larger differences in document mappings. However, Jaccard similarity is still important to consider, especially for datasets comprised of documents that are very semantically similar.
5.2 Cohesiveness and Distinctiveness
Pacheco et al. (2023) Fang et al. (2023)
Intra-Theme Inter-Theme Intra-Theme Inter-Theme
Sync $\bf 0.51(0.08)^{*}$ $0.41(0.04)^{*}$ $0.52(0.10)$ $0.40(0.04)$
Async $0.45(0.10)^{*}$ $\bf 0.34(0.11)^{*}$ $0.52(0.10)$ $0.40(0.04)$
Table 2: Group Avg. Similarity within Annotator Groups (All data points). Themes are considered to be more cohesive if intra-theme similarity is high, and more distinctive if inter-theme similarity is low. Statistically significant using a two-sample unpaired t-test with $p<0.05$.
A second dimension for determining the "systematicity, clarity, and transparency" of annotation outcomes is by evaluating the similarities and differences between themes within the same code-book (Cornish et al., 2014). We propose two metrics, cohesiveness and distinctiveness, to represent these similarities and differences respectively. Whereas consistency measures the systematicity and clarity of themes across annotators, cohesiveness and distinctiveness measure these qualities for themes by the same annotator. The distinctiveness of a theme is determined by comparing it to other themes created by the same annotator. On the other hand, the cohesiveness of a single theme is determined by its semantic homogeneity. A theme is said to be cohesive if its documents are similar to each other (measured by intra-theme similarity) and distinctive if it is dissimilar from documents in other themes within the same code-book (measured by inter-theme similarity). Intuitively, the purpose of grouping documents by theme is to create abstract representations of a dataset, where each theme represents a distinct facet of the data. If themes are not cohesive and distinctive, then it becomes hard to tell which theme a given document should belong to and the code-book falls apart.
To evaluate these metrics for a single annotator (or annotator group), we calculate both the intra-theme similarity and the inter-theme similarity for theme in the code-book. Intra-theme similarity is calculated by taking the average of pair-wise similarity between all documents of the same theme. Inter-theme similarity for a given theme is calculated by taking the average pair-wise similarity of documents in that theme with documents in all other themes by the same annotator.
A confounding factor in these measures is that both systems provide broad coverage of documents such that even distantly related documents may be assigned to a theme. In order to more accurately represent the cohesiveness and distinctiveness of themes in each experiment, we perform the same calculations on a subset comprised of only the top 25% of documents most closely related to each theme. For the Pacheco et al. (2023) system, this top quartile is selected using the distance from the centroid. For the Fang et al. (2023) system, the distribution of theme weights better reflect the assignments provided by the topic model and are used instead.
Table 2 shows results for the whole dataset and Table 3 show the same metrics for the subset of the documents closest to each theme. Overall, we find that the intra-theme similarities are always higher than inter-theme similarities, which means that themes are at least moderately cohesive and distinctive across the board. For the Pacheco et al. (2023) experiments, we find that the difference between both types of metrics remains stable across collaboration modalities, and that similarities are higher in the synchronous setting. Our findings hint that themes may be more cohesive but not more distinctive in the synchronous setting.
Pacheco et al. (2023) Fang et al. (2023)
Intra-Theme Inter-Theme Intra-Theme Inter-theme
Sync $\bf 0.70(0.08)^{*}$ $0.52(0.07)^{*}$ $0.56(0.11)$ $0.39(0.05)$
Async $0.64(0.09)^{*}$ $\bf 0.46(0.13)^{*}$ $0.56(0.10)$ $0.39(0.05)$
Table 3: Group Avg. Similarity within Annotator Groups (25% most related). Themes are considered to be more cohesive if intra-theme similarity is high and more distinctive if inter-theme similarity is low. Statistically significant using a two-sample unpaired t-test with $p<0.05$.
The results in the Fang et al. (2023) experiments are nearly identical across collaboration modalities. Narrowing down to the top 25% of closest documents increases the difference between the two metrics, lining up with the intuitive notion that doing so increases both cohesiveness and distinctiveness. When comparing the two systems, we find that the themes coded using the Pacheco et al. (2023) system are significantly more cohesive when looking at the quartile subset, which matches our findings in Section 5.3.
5.3 Correctness
While the previous evaluations look at the code-book as a hole, we want to also evaluate how correct our document to theme assignments are. In order to verify this, we conducted a post-hoc analysis of how accurate theme assignments are on a held-out set of randomly sampled tweets, compared across synchronous and asynchronous settings. The data was sampled after the annotation process was over. This meant that for both systems, annotators had deemed the theme discovery and refinement to have reached saturation.
Following Section 5.2, we perform one analysis of correctness that considers the set of all tweets, as well as a second analysis considers only the top $25\%$ of tweets most closely related to each theme. For each analysis, we sampled $200$ tweets randomly and had two of the authors of the paper manually annotate whether the theme assigned to a tweet by the system makes sense or not. We did this for each system and collaboration setting (synchronous vs. asynchronous), with the exception of the Pacheco et al. (2023) synchronous setting, for which we report the results presented in the original paper. We made sure to follow the same sampling procedure so that results were comparable.
Results are presented in Tab. 4. We observe that for both systems, the accuracy yielded from synchronous annotations is higher than their asynchronous counter parts. We also see that as per our expectations, the top $25^{th}$ percentile has a higher accuracy than group average of all tweets.
However, it is interesting to observe the comparatively poor accuracy of the topic modeling system Fang et al. (2023) as compared to the theme discovery system Pacheco et al. (2023). We attribute this difference to the fact that topic modeling systems do not "learn" to classify themes, but rather learn the probability that a given word belongs to a particular topic. During evaluation, we found that the main problem arises when salient words are used in atypical contexts, or when trying to bifurcate terms that are used in opposing themes.
As an example, take the following headline regarding the COVID-19 vaccine mandate.
Indiana University students must comply with the school’s COVID-19 vaccine mandate for the fall semester after a federal judge refused on Monday to block the school’s inoculation requirement, one of the first rulings of its kind during the pandemic
The topic modeling system gave the highest wright to the topic Anger over mask mandates. But really, this is regarding vaccines. The term mandate biases the model to rank a topic about masks (the context that typically appears around the word "mandate") higher than a topic related to the COVID-19 vaccine, which this headline is about.
Our findings regarding the correctness can be summarized by acknowledging that synchronous teams of annotators achieve significantly higher assignment accuracy than asynchronous teams. We also find that topic modeling systems achieve un-competitive assignment accuracy versus theme discovery systems. Lastly, qualitative researchers may want to choose theme discovery tools if they want higher accuracy in document to theme assignments – but could opt for topic modeling systems if they want to put their focus on word usage patterns.
Pacheco et al. (2023) Fang et al. (2023)
All Top 25% All Top 25%
Sync ${}^{*}68\%$ ${}^{*}86\%$ $23\%$ $42\%$
Async $50\%$ $64\%$ $17\%$ $35\%$
Table 4: Correctness: Accuracy of Theme Assignment for Held-out Dataset. Results taken from Pacheco et al. (2023).
5.4 User Study
We carried out semi-structured interviews with all the participants in the study after they were done with their coding task. We focused on trying to understand attitudes towards the task and tools, particularly with respect to the synchronous and asynchronous modalities. We summarize our main findings below.
Synchronous annotators found working in teams to ease the annotation process.
Both teams of annotators working synchronously using the Fang et al. (2023) described their perceived advantage of working in teams – broader perspectives by inviting view points from different annotators helped them contextualize diverse tweets easily. The annotators also valued the ability to collaboratively resolve contested tweets through discussions. Particularly, it was easy to "break ties" quickly through conversation when it came to mapping which words belonged to a given a topic. The annotators were also quicker to reach saturation, and they expressed that they thought this was possible because consolidation was possible on-the-fly. One of the annotators noted that a pre-requisite for successful synchronous coding stemmed from the rapport their annotation team shared, as they are all members of the same lab and were able to convey their viewpoints without hesitation.
Asynchronous annotators were more likely to notice inadequacies of the annotation tool they used.
While most of the highlights from the synchronous annotators focused on the process of annotation, the asynchronous annotators directed their feedback towards the way the tools functioned. Since the bulk of the annotations were carried in isolation, difficulties navigating the tool were more salient to them than to their synchronous counterparts. The synchronous annotators also met with similar issues, but were quick to solve it collaboratively. Even though all annotators (in both settings and systems) started with the same initial clusters of tweets, the synchronous annotators were able to find the operation they need performed quicker through in-team collaboration.
Annotators noted the lack of control and agency that the topic modeling system had over the theme discovery system.
Annotators expressed loss of control and agency when using the topic modeling system Fang et al. (2023) over the theme discovery system Pacheco et al. (2023). One annotator noted that the controls to modify topics were not satisfactory – "the merge process did not offer the ideal amount of control and made it difficult to keep track of the theme groups".
Annotators felt that starting topics often negatively impacted their annotations.
Since topic models start by a sampled cluster of data, both synchronous and asynchronous annotators were dissatisfied with some of the ways in which tweets belonging to different or opposite themes were placed under the same topic. One of the annotators commented – "Many Anti-Vax and Pro-Vax standpoints use the same words/phrases in their tweets, which the Fang et al. (2023) model groups together despite the stark difference in message between the two.". Since the topics contain a large number of documents, it is only feasible for annotators to include / exclude a very small subset as being part of the topic. Some users felt frustrated at being unable to achieve the purity of topic that they hoped for, often stopping when they found that none of their changes were affecting the topics as they intended. However, one annotator did note that the benefit of using a topic modeling system was that there was some amount of clustering already done - as opposed to the Pacheco et al. (2023) system where annotations had to be done from scratch, posing to be a small hurdle in the beginning.
6 Conclusion and Future Work
In this paper, we examined NLP-assisted qualitative research tools in different collaboration modalities. We conducted annotations on a vast dataset using two popular interactive coding tools, Pacheco et al. (2023) and Fang et al. (2023). We then designed an evaluation framework that describes the quality of the induced themes and their resulting document assignments under synchronous and asynchronous collaboration. We note that in general, the collaboration modality is a significant factor in determining if the quality of a system’s output. This is particularly true for theme discovery systems like Pacheco et al. (2023), where users have a wider bandwidth of interaction that can benefit from group consensus. We also observe that solutions based on topic modeling, although popular in the data analysis literature, can struggle with inducing cohesive themes and highly accurate document-level annotations.
While this study focuses on collaboration modalities, there are numerous other variables that can affect a tool’s efficacy in supporting qualitative coding. We believe that the evaluation framework that we proposed can be repurposed and expanded to evaluate a wide range of interventions, such as the underlying NLP technology, the interactive interface, the expertise and life experience of the coders, and the type of data being annotated. Through this paper, we hope to inform future work that gives way to more robust measures of an NLP tool’s capabilities as it pertains to qualitative research under realistic settings.
7 Limitations
The study presented in this paper has three main limitations.
(1) While we selected two distinct tools to perform our analysis of synchronous vs. asynchronous settings, these tools are not representative of the full range of solutions that have been proposed to support qualitative coding. In particular, neither of these systems utilize large language models (LLMs), which have been at the forefront of recent developments in the field Dai et al. (2023); Gao et al. (2024), and which have garnered significant interest from qualitative researchers Hayes (2023); Roberts et al. (2024); Aubin Le Quéré et al. (2024).
(2) As prior work has shown Hoyle et al. (2021), automated metrics and human evaluation do not always correlate well for unsupervised or semi-supervised theme induction tasks like topic modeling. Although we did incorporate two types of human evaluation: a post-hoc manual verification (Sec 5.3) and a user study (Sec 5.4), both of these evaluations were limited in scope. This is primarily due to how expensive and time consuming it is to perform an exhaustive human evaluation. While we are confident in the findings of our study, future work should expand on the human analysis of these solutions to provide a more robust view.
(3) While we look at an important variable in qualitative research settings (collaboration modality), there are several other variables that can influence the outcome of NLP-assisted solutions (e.g., choice of tool, expertise and live experience of annotators, type of data being annotated, etc.). In addition to this, we did not explore the many different consolidation strategies that are often used to bring together the perspectives of asynchronous coders. We leave the explorations of these questions for future work.
8 Ethical Considerations
To the best of our knowledge, no code of ethics was violated during the development of this project. We used publicly available tools and datasets according to their licensing agreements. For our annotation experiments, we followed IRB protocol and did not retain any personally identifiable information.
All information needed to replicate our experiments is presented in the paper. We reported all experimental settings, as well as any pre-processing steps, learning configurations, hyper-parameters, and additional technical details. Due to space constraints, some of this information was relegated to the Appendix. In addition to this, we will make the results of the annotation experiment available to the community, as well as the code to produce all of our reported results***https://github.com/blast-cu/interactive-systems .
Appendix A Interactive System Details
Operations Description
Finding Partitions Experts can find partitions in the space of unassigned instances. We currently support the K-means Jin2010 and Hierarchical Density-Based Clustering mcinnes2017hdbscan algorithms.
Text-based Queries Experts can type any query in natural language and find instances that are close to the query in the embedding space.
Finding Similar Instances Experts have the ability to select each instance and find other examples that are close in the embedding space.
Listing Themes and Instances Experts can browse the current list of themes and their mapped instances. Instances are ranked in order of “goodness”, corresponding to the similarity in the embedding space to the theme representation. They can be listed from closest to most distant, or from most distant to closest.
Visualizing Local Explanations Experts can visualize aggregated statistics and explanations for each of the themes. To obtain these explanations, we aggregate all instances that have been identified as being associated with a theme. Explanations include wordclouds, frequent entities and their sentiments, and graphs of concept distributions.
Visualizing Global Explanations Experts can visualize aggregated statistics and explanations for the global state of the system. To do this, we aggregate all instances in the database. Explanations include theme distribution, coverage statistics, and t-sne plots vanDerMaaten2008.
(a) Exploratory Operations
Operations Description
Adding, Editing and Removing Themes Experts can create, edit, and remove themes. The only requirement for creating a new theme is to give it a unique name. Similarly, themes can be edited or removed at any point. If any instances are assigned to a theme being removed, they will be moved to the space of unassigned instances.
Adding and Removing Examples Experts can assign “good” and “bad” examples to existing themes. Good examples are instances that characterize the named theme. Bad examples are instances that could have similar wording to a good example, but that have different meaning. Experts can add examples in two ways: they can mark mapped instances as “good” or “bad”, or they can directly contribute example phrases.
Adding or Correcting Concepts We allow users to upload additional observed or predicted concepts for each textual instance. For instances and phrases added as “good” and “bad” examples, we allow users to add or edit the values of these concepts. The intuition behind this operation is to collect additional information for learning to map instances to themes.
(b) Intervention Operations
Table 5: Interactive Operations for the Pacheco et al. (2023) System